Grade 11Organic Chemistry

log[H+]=4.15 then how what is the value of [H+] and how to calculate that? please explain your answer quickly.

Reddy Rakesh
13 Years agoGrade 11

4 Answers

AAYUSH KESHAVA
ApprovedApproved Tutor Answer13 Years ago

log[H+]=4.15 ie log10[H+]=4.15

taking antilog both sides;

antilog log10[H+]=antilog 4.15

so,

[H+]=10^(4.15)

using log table we get 

[H+]=10^(4.15)=14125.375

it is important to note anti log is taken to remove log . base should be appropriate while using logarithemic functions.

Tanmay Warwadkar
13 Years ago

Take log of 4.15.... You will get the answer....

Har Simrat Singh
13 Years ago

Believing its log to the base 10 using Loga(x)= y then x=a^y

hence [H] = 10^4.15 this thing is called antilog 

if allowed a log book look up the value of .15 in antilog portion ,say it is 1234 then [H] = 1.234 x 10^4

mannam kiran kumar
ApprovedApproved Tutor Answer13 Years ago

dear friend ,

                          if you got any type antilog sums one tip you have to learn that is

                                 antilog (x) = 10^x

                this is useful for any problem ok

                  and solved this type of questions. good luck